OpenStudy (anonymous):

Quadratic equations that can be solved using the quadratic formula can also be solved by factoring. Question 1 options: 1) always 2) sometimes 3) never

OpenStudy (chaise):

sometimes

OpenStudy (anonymous):

always just not with integers

myininaya (myininaya):

this question is stupid

OpenStudy (anonymous):

i'll say. some math teacher problem

myininaya (myininaya):

i mean you arent stupid for asking it marc this is a stupid question to assign

OpenStudy (chaise):

Yeah, silly question. It can be factored if you don't use intergers, which is a somewhat difficult and lengthy process.

OpenStudy (anonymous):

I really dont understand that last problem chaise..Solve 3x2 + 4x = 2. Question 3 options: 1) 2 (+or-) √10 all over 6 2) -2 (+or-) √10 all over 3 3) -2 (+or-) √10 all over 3 4) -4 (+or-) √10 all over 3 can anyone help

OpenStudy (anonymous):

right. factor theorem says if r is a root of a polynomial p(x) then p(x) can be factored as \[p(x)-(x-r)q(x)\]

OpenStudy (anonymous):

use quadratic formula after writing as \[3x^2+4x-2=0\]

OpenStudy (anonymous):

i.e. use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=3,b=4,c=-2\]

OpenStudy (anonymous):

get \[x=\frac{-3\pm\sqrt{10}}{3}\] but it requires writing \[\sqrt{16+24}=\sqrt{40}=\sqrt{4\times 10}=2\sqrt{10}\] and then canceling a 2 from top and bottom

OpenStudy (anonymous):

that is why you get the answer i wrote.

OpenStudy (anonymous):

The choice options are -2 and -4 \[\pm √10 / 3\]

myininaya (myininaya):

\[x=\frac{-4 \pm \sqrt{4^2-4(3)(-2)}}{2(3)}=\frac{-4 \pm \sqrt{40}}{6}\] \[=\frac{-4 \pm 2\sqrt{10}}{6}=\frac{-2 \pm \sqrt{10}}{3}\]

myininaya (myininaya):

i think satellite meant to write a -2 not a -3 above anyways the answer is what i have

OpenStudy (anonymous):

Thats what i thought